40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Message ID: 165     Entry time: Tue Sep 3 17:03:25 2013
Author: Koji 
Type: General 
Category: General 
Subject: H1 OMC gluing completed 

[Koji Jeff]

H1 OMC All Gluing completed

5 Glue H1 beam dumps (UV)

4 glass wire brackets glued on the H1 topside (UV) SN: #9/10/11/12

6 Invar blocks glued on the H1 topside (EP30) SN: #13/14/15/16/18/19

ELOG V3.1.3-